An air conditioning system can circulate 450 cubic feet of air per minute. How many cubic yards of air can it circulate per minute
Found 2 solutions by MathLover1, Alan3354:
Answer by MathLover1(20849)
Answer by Alan3354(69443)

An air conditioning system can circulate 450 cubic feet of air per minute. How many cubic yards of air can it circulate per minute
=========
1 yd = 3 ft
1 sq yd = 3^2 sq ft = 9 sq ft
1 cu yd = 3^3 cf = 27 cf
---> 450/27 cyd